OpenStudy (anonymous):

If x = -3 then find the matching value for y. 5x-3y=8 ANSWER please.

OpenStudy (diyadiya):

Substitute the value x in the equation 5x-3y=8 5(-3) - 3y = 8 can you do it now?

OpenStudy (btaylor):

plug -3 in for x in the equation. It turns into: 5(-3) - 3y = 8 -15 - 3y = 8 -3y = 23 y = -23/3

OpenStudy (anonymous):

would the answer be -7.6?

OpenStudy (anonymous):

yup but its 7.666666... so as we write 7.667